OpenStudy (anonymous):

i reallyyyy need help!!!! Please someone find it in ur heart to help me! lol

OpenStudy (anonymous):

For all 5 problems in this assignment, use the following functions: f(x) = x + 5 g(x) = x2 -4x + 5 h(x) = 3x2 - 2x + 5

OpenStudy (anonymous):

Find (f+g)(x)

OpenStudy (anonymous):

distribute the x (f+g)(x) = f(x)+g(x) can you do f(x)+g(x) ?

OpenStudy (anonymous):

but wat is x?

OpenStudy (anonymous):

x is a variable it is not asking you to solve.

OpenStudy (anonymous):

it is asking you to add the 2 functions f(x) = x + 5 g(x) = x2 -4x + 5 f(x)+g(x)=( x + 5)+( x2 -4x + 5) x^2-3x+10 add like terms

OpenStudy (anonymous):

hmm ohhhhh ohk hold on

OpenStudy (anonymous):

2x+5?

OpenStudy (anonymous):

i didn't explain right maybe but x^2-3x+10 is the answer

OpenStudy (anonymous):

here is a visual way

OpenStudy (anonymous):